OpenStudy (anonymous):

If f '(2) = 3.1 and g '(2) = 7.3, then the graph of f(x) + g(x) has slope 10.4 at x = 2

OpenStudy (anonymous):

yes is that the question?

jimthompson5910 (jim_thompson5910):

h(x) = f(x) + g(x) h'(x) = f'(x) + g'(x) h'(2) = f'(2) + g'(2) h'(2) = 3.1 + 7.3 h'(2) = 10.4

OpenStudy (anonymous):

what about this one? also true or false question If y satisfies the equation y2 + xy − 1 = 0, then dy dx exists everywhere.

jimthompson5910 (jim_thompson5910):

what did you get

jimthompson5910 (jim_thompson5910):

a graph will be really handy since you can visually confirm or deny the statement

OpenStudy (anonymous):

no, that is all the information given...

jimthompson5910 (jim_thompson5910):

I know, but you can still graph it

jimthompson5910 (jim_thompson5910):

here's what the graph looks like

jimthompson5910 (jim_thompson5910):

notice how it's continuous on each branch

jimthompson5910 (jim_thompson5910):

and how there aren't any sharp corners or cusps

OpenStudy (anonymous):

did you use a calculator to graph it and if so how do you graph an equation with two variables in it?

jimthompson5910 (jim_thompson5910):

well you would have to solve for y first if you were to do it on a graphing calculator I used geogebra to graph it (since it can handle implicit plots)

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

yw